# Interpretability Needs a New Paradigm

### v1

@@ 4.1 The case against the intrinsic paradigm
Another challenge with the intrinsic paradigm is that its models are often not completely interpretable because only a part of the model is aritectually constrained to be interpretable. The rest of the neural network, still use black-box components (e.g. Dense layer, Recurrent layer, etc.) which are not interpretable. As such, the intrinsic promise should not be taken at face value (Jacovi & Goldberg, 2020).
An example of this, is classic attention-based models (Bahdanau et al., 2015; Jain & Wallace, 2019). Attention itself is interpretable, as it’s a weighted sum, and explains the importance of each intermediate representation. However, attention is often used as token-importance. This does not work, as the intermediate representations are produced by a black-box recurrent neural network (e.g., LSTM Hochreiter & Schmidhuber 1997) which can mix or move the relationship between tokens and the intermediate representations. Therefore, the attention scores do not necessary represent token-importance (Bastings & Filippova, 2020).
Likewise, Neural Modular Networks produce an executable problem composed of sub-networks, such as find-max-num(filter(find())), which is interpretable (Fashandi, 2023; Andreas et al., 2016; Gupta et al., 2020). However, each sub-networks (find-max-num, filter, find) is itself a black-box model with little guarantee that it operates as intended (Amer & Maul, 2019; Subramanian et al., 2020; Lyu et al., 2022).
@@ 4.2 The case against the post-hoc paradigm
Most notable is perhaps post-hoc importance measure (IM) explanations, where the explanation indicates which input features are the most important for making a prediction. The pursuit of such explanations have produced countless papers (Binder et al., 2016; Ribeiro et al., 2016; Li et al., 2016; Shrikumar et al., 2017; Smilkov et al., 2017; Sundararajan et al., 2017; Ahern et al., 2019; Thorne et al., 2019; ElShawi et al., 2019; Sangroya et al., 2020).
However, repeatedly, the faithfulness of these IM explanations is criticized (Adebayo et al., 2018, 2021; Kindermans et al., 2019; Hooker et al., 2019; Slack et al., 2020; Yeh et al., 2019). For example, there is great disagreement between alleged faithful IMs, which is hard to reconcile (Jain & Wallace, 2019; Krishna et al., 2022). Other works show their faithfulness is both task and model-dependent and thus don’t provide the generality that the post-hoc paradigm desires (Bastings et al., 2022; Madsen et al., 2022a). Finally, theoretical works suggest that IMs are subject to a no free lunch thorem (Han et al., 2022), or it may be impossible to provide faithful post-hoc IMs (Bilodeau et al., 2024).
@@
Unfortunately, like post-hoc importance measures, there are many reasons to be highly skeptical (Belinkov, 2021). For example, using an untrained model or a randomized dataset shows an equally high correlation with linguistic properties, compared with training a regular model (Zhang & Bowman, 2018; Hewitt & Liang, 2019). These discoveries have put the entire methodology into question, although there is work trying to adapt to these new critics (Voita & Titov, 2020).
4.3 Conclusion
@@
Although both the intrinsic and post-hoc paradigms have significant issues, parts of their underlying beliefs have merit. The intrinsic paradigm believes that we can’t expect models that were not designed to be explained to be explained, while post-hoc believes black-box models tend to be more general purpose and while providing high predictive performance. These beliefs have merit, and it’s worth considering how to incorporate their spirit into new paradigms.
5 Possible new paradigms
@@
[…]
Given the shortcomings of the intrinsic and post-hoc paradigms, we are now in a situation where it’s meaningful to look beyond these current paradigms. To prove that developing alternative paradigms is possible and to highlight some new directions of research, this section discusses new directions for interpretability, frames them as paradigms, and summarizes them in Table 2. However, these new paradigms are not the only directions one could take for interpretability, and in general, we will encourage researchers to consider alternatives that have not yet been explored.
5.1 Inherently Faithfulness Measurable Models
The faithfulness of an explanation can be extremely difficult to measure because the true explanation is inaccessible to us (Jacovi & Goldberg, 2020). Therefore, faithfulness metrics often have to make use of a proxy measure.
For example, the erasure metric for importance measures says: if information (pixels, tokens, etc.) is truly important, then when removing it the prediction should change significantly (Samek et al., 2017). However, this has various limitations, because removing information causes out-of-distribution issues, which other works have solved by repeatedly retraining the model on datasets with allegedly important information removed (Hooker et al., 2019). Unfortunately, retraining is computationally expensive, and measuring on a retrained model means faithfulness is no longer measured on the original model. Other faithfulness metrics have similar limitations (Madsen et al., 2024b).
Therefore, rather than considering a faithfulness metric as something that exists externally of the model, the idea with Inherently Faithfulness Measurable Models (FMMs) is a model that inherently provides the means to cheaply and accurately measure faithfulness (Madsen et al., 2024b). Then, because the faithfulness measure is cheap and accurate, it’s possible to optimize an explanation to maximize this faithfulness metric (Zhou & Shah, 2023).
This is different from how inherently explainable models in the intrinsic paradigm which are directly designed to be explained because their architectural constraints produce the explanation (Jain & Wallace, 2019; Rudin, 2019). The central idea in FFMs is the model does not produce an explanation directly but enables a faithfulness metric, thus making FFMs indirectly designed to be explainable (Madsen et al., 2024b).
In essence, Madsen et al. (2024b) reformulates the interpretability problem from ‘inherently explainable” to “inherently measurable” and show that this reformulation is much easier to design for, as it does not require architectural constraints. The concrete demonstration that Madsen et al. (2024b) provide, uses a RoBERTa model (Liu et al., 2019) where the loss function is modified such that removing tokens becomes in-distribution, thereby enabling the erasure metric (Samek et al., 2017) to be applied directly without retraining. This achieves state-of-the-art faithfulness metrics without affecting performance or inference time.
Because the work done by Madsen et al. (2024b) only applies to masked language models. Thus, the immediate next step is to extend these ideas to causal language models which may be a challenging task as their idea depends on mask-tokens. Additionally, new work may explore more ways of optimizing explanations given a cheap faithfulness metric (Zhou & Shah, 2023).
Beyond these immediate applications, the principle of enabling models to be faithfulness measurable can be applied to other explanations. For example, concepts (Kim et al., 2018) and influential examples (Yeh et al., 2018; Pruthi et al., 2020; Koh & Liang, 2017) may be good candidates for this. Both the general idea and the specific approach by Madsen et al. (2024b), can also be applied to other domains such as computer vision.
To summarize, while previous interpretability works have explained first and then measure faithfulness later, this paradigm does the opposite; it designs for faithfulness measures first and then explains. This reformulation of the interpretability problem is easier to design for, thus satifing both the black-box and designed-to-be-explained desirables.
5.2 Learn to faithfully explain
Rather than optimizing the predictive model to be faithfulness measurable, a more long-standing idea in interpretability has been to optimize the predictive model such that an explanation method applied to that model produces faithful explanations (Yoon et al., 2019; Chen et al., 2018). This also does not require the architectural constraints that the intrinsic paradigm applies, as the explanation comes from an external explanation method, not the architectural design.
The explanation method can be similar or even identical to those from the post-hoc paradigm. However, because the model is optimized to enable these explanations to be faithful, it’s not post-hoc, and there are more reasons to think that the explanations should be faithful.
Early work on this used an explanation model, which will produce an explanation of the prediction model. The explanation and prediction models are then trained jointly to produce faithful explanations and high-accuracy predictions (Yoon et al., 2019; Chen et al., 2018). This direction has been called joint amortized explanation methods (JAMs). However, Jethani et al. (2021) point out that the explanation model often learns to encode the prediction, which means the explanation model becomes part of the black-box problem rather than the solution.
Jethani et al. (2021) solve this by using a disjoint setup, where the explanation model can’t encode the prediction, a setup that following works have adapted (Jethani et al., 2022; Covert et al., 2022). However, this still presents a problem as it assumes the explanation model is 100% robust. For example, an out-of-distribution observation could break the explanation model and produce unfaithful explanations. This is particularly relevant, as it is often the cases where models break that are desirable to explain.
Instead, a better direction is likely to use a simple explanation algorithm that is easy to understand rather than an explanation model. Bhalla et al. (2023) do this by jointly training a predictive model with an explanation algorithm, where explanations indicate unnecessary features and the model learns to support sparse features.
Regardless of the specific approach used to produce the explanation, the optimization criteria used to make the predictive model converge, such that the explanations are faithful, is either a relaxation of a faithfulness metric or an underlying criterion necessary to produce faithful explanations. For example, masking inputs should be in-distribution (Covert et al., 2022; Bhalla et al., 2023). Therefore, this field needs to work on how to ensure and validate that the explanations are actually faithful and that the model properties that make explanations faithful also hold for out-of-distribution data.
Another direction in this paradigm is the idea that regularization can produce more faithful explanations. For example, a more locally-linear model may be more suitable for a gradient-based explanation (Ross & Doshi-Velez, 2018; Srinivas et al., 2022; Geirhos et al., 2023), or training a computer vision model with added input-noise reduces noise in importance measure explanations because the model becomes less sensitive to one-pixel variations (Bansal et al., 2020). As of now, these observations are mostly hypothetical or anecdotal, and we advocate for a systematic study from an interpretability perspective.
5.3 Self-explaining models
Rather than using external algorithms or models to produce explanations, Elton (2020) proposes that models should explain themselves, meaning they become self-explaining. Because there are now two black-box modalities (the self-explaining and the normal prediction modality), this can be quite dangerous, and there must be solid evidence that the explanations are faithful for this approach to be valid.
Despite the immediate dangers with a self-explaining model, the best tool to explain a black-box model may be the model itself. In particular, unlike the JAM approach (mentioned in Section 5.2), a self-explaining model has full access to its own internal weights and activations. The explanations can also be much more informative. This particularly holds for instruction-tuned large-language models (LLMs) (OpenAI, 2023; Jiang et al., 2023; Meta, 2023), which are allegedly able to explain themselves in great detail and very convincingly (Chen et al., 2023; Agarwal et al., 2024).
When LLMs explain themselves, this is referred to as self-explanations. Importantly, self-explanations must relate to the model’s reasoning logic, not just the world or abstract concepts (e.g., mathematics), which is how most Chain-of-Thought (CoT) explanations are evaluated (Chu et al., 2023). As humans don’t know the model’s logic, evaluating self-explanations is difficult (Huang et al., 2023).
Chen et al. (2023) show although self-explanations are very convincing to humans (called plausibility), they do not help humans to guess model predictions of new observations (called simulatability) and, therefore, can be dangerous. They argue that “naively optimizing human approvals (e.g., RLHF)” does not improve self-explanations’ simulatability.
Simulatability is important for measuring the usefulness of explanations. However, it is difficult to ensure that the human test subjects make their guesses based on the model’s self-explanation and not their world knowledge. Therefore, measuring faithfulness is equally important, as this evaluates explanations independently of humans.
To measure faithfulness in LLMs, one can use self-consistency checks (Madsen et al., 2024a; Huang et al., 2023; Lanham et al., 2023). For example, Lanham et al. (2023) considers CoT explanations for mathematical reasoning, edits facts in the CoT (e.g., from “2+2=4” to “2+3=5”), and then checks that the prediction changes based on these edits. Another example is Madsen et al. (2024a), who evaluate counterfactual and importance measure explanations by asking a model to change the input such it gives a different prediction; this can be verified by re-evaluating with the counterfactual input. Both Lanham et al. (2023) and Madsen et al. (2024a) find that the faithfulness of LLMs’ self-explanations is model and task-dependent.
Unfortunately, self-consistency checks are impossible for most explanations (Madsen et al., 2024a; Lanham et al., 2023). An alternative may be evaluating the model’s ability to model itself (called self-modeling). The hypothesis is that self-modeling is a necessary condition for general faithful self-explanations (Madsen et al., 2024a).
Anthropic Team (2022) evaluate self-modeling capabilities (not in a self-explanation context) by asking the LLM if it knows the answer to a question and comparing it with simply asking the LLM to answer the question. They find that LLMs cannot perform this task without additional fine-tuning. Even with fine-tuning, the LLM cannot generalize to other domains. As such, developing general self-modeling capabilities could be a first step towards general self-explanation. Such work should also show that self-modeling capabilities make concrete explanations faithful too, using self-consistency checks (Madsen et al., 2024a).
5.4 Comparison
@@
All three mentioned paradigms work with what-would-be black-box models. However, their idea is to optimize these models to guarantee a faithful explanation. They optimize for either faithfulness measurability, explanation faithfulness, or self-model/self-explanation capabilities.
In conclusion, they are different from the post-hoc paradigm because interpretability is considered during optimization, and they are different from the intrinsic paradigm because they are not architecturally constrained. They are also mutually distinct, each taking a different perspective on achieving faithful interpretability via optimization.
At first, this sounds like a natural and simple idea. However, recall that we can’t know what a faithful explanation is or what it looks like, as the model’s internal logic is inaccessible to us. Achieving interpretability via optimization is thus extremely challenging and involves proxies, assumptions, etc., which is how they differ.
@@ 6 Limitations
This concern is orthogonal to the faithfulness concern and is therefore not discussed in this paper. However, readers are encouraged to study recent works like Schut et al. (2023); Kim (2022), which proposes the new idea that it is not enough to frame explanations in terms that humans already understand. We should also develop new language and mental abstractions for humans to understand machines.
@@ 7 Conclusion
Although some evidence exists for the new paradigms that were presented in Section 5, these are, first and foremost, just ideas. It’s only in retrospect that we can truly know if one paradigm results in meaningful progress in the field. It is also entirely possible that neither of these ideas is what moves the interpretability field forward.
@@
Historically, a common tactic in post-hoc works was convincing visualization of explanations that aligned with our intuitions (Olah et al., 2017; Yosinski et al., 2015; Nguyen et al., 2016). However, such visualizations are empty arguments, as humans can’t know what a true explanation looks like (Geirhos et al., 2023). Likewise, intrinsic works have made seemingly strong theoretical arguments for why their methods were faithful, but these arguments failed to capture the whole model. Even the new learn-to-faithfully-explain paradigm have already shown sharp corners, where the explainer models unintentionally encode the prediction and are therefore unfaithful (Jethani et al., 2021).
@@
Finally, while these new paradigms are promising, it’s unlikely they will completely erase the current paradigms. We still teach both the particle and wave perspectives in physics. Most scientists don’t worry about whether there are true statements in math that cannot be proven.

### v2

@@ 3.1.1 The intrinsic paradigm
Figure 1: Abstract diagram of the intrinsic paradigm, where the model is architecturally constrained, such that the constraint itself is the explanation. In cases of Decision Trees the entire model is constrained, but often (e.g. Prototype Networks or Attention) only part of the model is constrained.
@@ 3.1.2 The post-hoc paradigm
Figure 2: Abstract diagram of the post-hoc paradigm, where a post-hoc method is used to explain a black-box model. The post-hoc method is usually an algorithm, like the gradient w.r.t. the input, but it can also be an auxiliary model.
@@ 4.1 The case against the intrinsic paradigm
Another challenge with the intrinsic paradigm is that its models are often not completely interpretable because only a part of the model is architecturally constrained to be interpretable. The rest of the neural network, still use black-box components (e.g. Dense layer, Recurrent layer, etc.) which are not interpretable. As such, the intrinsic promise should not be taken at face value (Jacovi & Goldberg, 2020).
An example of this is classic attention-based models (Bahdanau et al., 2015; Jain & Wallace, 2019). Attention itself is interpretable, as it’s a weighted sum, and explains the importance of each intermediate representation. However, attention is often used as token-importance. This does not work, as the intermediate representations are produced by a black-box recurrent neural network (e.g. LSTM Hochreiter & Schmidhuber 1997) which can mix or move the relationship between tokens and the intermediate representations. Therefore, the attention scores do not necessarily represent token-importance (Bastings & Filippova, 2020).
Likewise, Neural Modular Networks produce an executable problem composed of sub-networks, such as find-max-num(filter(find())), which is interpretable.(Fashandi, 2023; Andreas et al., 2016; Gupta et al., 2020). However, each sub-networks (find-max-num, filter, find) is itself a black-box model with little guarantee that it operates as intended (Amer & Maul, 2019; Subramanian et al., 2020; Lyu et al., 2024).
@@ 4.2 The case against the post-hoc paradigm
Most notable is perhaps post-hoc importance measure (IM) explanations, where the explanation indicates which input features are the most important for making a prediction. The pursuit of such explanations has produced countless papers (Binder et al., 2016; Ribeiro et al., 2016; Li et al., 2016; Shrikumar et al., 2017; Smilkov et al., 2017; Sundararajan et al., 2017; Ahern et al., 2019; Thorne et al., 2019; ElShawi et al., 2019; Sangroya et al., 2020).
However, repeatedly, the faithfulness of these IM explanations is criticized (Adebayo et al., 2018, 2021; Kindermans et al., 2019; Hooker et al., 2019; Slack et al., 2020; Yeh et al., 2019). For example, there is great disagreement between alleged faithful IMs, which is hard to reconcile (Jain & Wallace, 2019; Krishna et al., 2022). Other works show their faithfulness is both task and model-dependent and thus don’t provide the generality that the post-hoc paradigm desires (Bastings et al., 2022; Madsen et al., 2022a). Finally, theoretical works suggest that IMs are subject to a no free lunch theorem (Han et al., 2022), or it may be impossible to provide faithful post-hoc IMs (Bilodeau et al., 2024).
@@
Unfortunately, like post-hoc importance measures, there are many reasons to be highly skeptical (Belinkov, 2021). For example, using an untrained model or a randomized dataset shows an equally high correlation with linguistic properties, compared with training a regular model (Zhang & Bowman, 2018; Hewitt & Liang, 2019). These discoveries have put the entire methodology into question, although there is work trying to adapt to these new critiques (Voita & Titov, 2020).
4.3 Summary
@@
[…]
@@
5 Are new paradigms possible?
Although both the intrinsic and post-hoc paradigms have significant issues, parts of their underlying beliefs have merit. The intrinsic paradigm believes that we can’t expect models that were not designed to be explained, to be explained, while post-hoc believes black-box models tend to be more general purpose while providing high predictive performance. These beliefs have merit, and it’s worth considering how to incorporate their spirit into new paradigms.
It can seem unlikely that such a paradigm can even exist. However, there is already some work that satisfies these desirables. In particular, we have identified 3 alternative paradigms, summarized in Table 2. These directions are all fairly new and unfortunately have not received much focus, likely due to favoritism towards existing paradigms (Kuhn, 1996).
@@ 5 Are new paradigms possible?
All three mentioned paradigms work with what-would-be black-box models. However, their idea is to optimize these models, such that they are designed to be explained. How they differ, is in their exact formulation of this approach.
It’s important to note, that it’s only with hindsight we can truly know if a new idea will become the next major paradigm, and it may be a fourth unknown idea that will become the next major paradigm. As such, the main purpose of this section is not to promote the next paradigm but rather to establish that it is possible to develop new interpretability paradigms.
5.1 The learn-to-faithfully explain paradigm
This paradigm is the most direct application of the optimization idea. An existing explanation algorithm (Bhalla et al., 2023) or model is used (Yoon et al., 2019; Chen et al., 2018), and the predictive model is then optimized to maximize both the predictive performance and the faithfulness.
Importantly, this approach does not require the architectural constraints that the intrinsic paradigm applies, as the explanation comes from an external explanation method, not the architectural design. The explanation method can be similar or even identical to those from the post-hoc paradigm. However, because the model is optimized to enable these explanations to be faithful, it’s not post-hoc, and there are more reasons to think that the explanations should be faithful.
Figure 3: Abstract diagram of the learn-to-faithfully explain paradigm. In most cases, this paradigm works by generating an explanation from the input, using either a model or an algorithm, this explanation is then fed into the predictive model, which has been optimized to respect the explanation.
Early work on this jointly trains an explanation model and a prediction model (Yoon et al., 2019; Chen et al., 2018). This direction has been called joint amortized explanation methods (JAMs). However, Jethani et al. (2021) point out that the explanation model often learns to encode the prediction, which means the explanation model becomes part of the black-box problem rather than the solution. A solution can be to use a disjoint setup (Jethani et al., 2021), where the explanation model can’t encode the prediction, a setup that following works have adapted (Jethani et al., 2022; Covert et al., 2022). However, the explanation model may still output unfaithful explanations for out-of-distribution inputs. An alternative is to produce the explanation algorithmically (Bhalla et al., 2023), for example by having an explanation algorithm remove unnecessary features, and the prediction model learns to support sparse features.
Regardless of the specific approach used to produce the explanation, the challenges are formalizing the faithfulness objective correctly such that the optimization works as intended, ensuring that the explanations are truly faithful and that the model properties that make explanations faithful also hold for out-of-distribution data (Covert et al., 2022; Bhalla et al., 2023).
5.2 The faithfulness measurable model paradigm
This paradigm integrates measuring the faithfulness of an explanation into the model design, such faithfulness can be easily measured without requiring architectural constraints. This can be of a huge advantage, as measuring faithfulness is often extremely challenging (Jacovi & Goldberg, 2020). Importantly, because faithfulness is easy to measure by design, it’s possible to identify the explanation that maximizes faithfulness using optimization algorithms (Zhou & Shah, 2023), which makes the model indirectly intrinsically explainable (Madsen et al., 2024b; Hase et al., 2021; Vafa et al., 2021). In essence, this paradigm reformulates the intrinsic paradigm from ‘inherently explainable” to “inherently measurable”.
Figure 4: Abstract diagram of the faithfulness measurable model paradigm. In this paradigm, the predictive model can also measure how faithful a given explanation is. The explanation can thus be produced by optimizing an initial (maybe random) explanation towards maximal faithfulness.
Madsen et al. (2024b) and Hase et al. (2021) show that this idea can be achieved using simple data argumentation, and there is no need for architectural constraints. The central idea is to use the erasure metric (Samek et al., 2017) to measure faithfulness of importance measures. The erasure metric says: if information (pixels, tokens, etc.) is truly important, then when removing it the prediction should change significantly. The common challenge is that removing information causes out-of-distribution issues (Hooker et al., 2019; Madsen et al., 2022a). However, by using data argumentation during training, it’s possible to extend the model to support the partial inputs created by the erasure metric. Importantly this can be achieved without architectural constraints, thus it remains possible to use general-purpose models such as RoBERTa (Madsen et al., 2024b) and GPT-2 (Vafa et al., 2021).
The challenge in this paradigm is about coming up with a way to integrate the faithfulness metric in the model, while ensuring there is no performance impact and that the model operates in-distribution (Madsen et al., 2024b). Additionally, developing efficient optimization procedures for optimizing explanations is difficult, due to the discrete nature of many explanations (Hase et al., 2021; Zhou & Shah, 2023).
5.3 The self-explaining model paradigm
Rather than using external algorithms or models to produce explanations, Elton (2020) proposes in this paradigm that models should explain themselves, meaning they become self-explaining. The most common implementation of this idea is instruction-tuned large language models (e.g., ChatGPT, Gemini, etc.) (OpenAI, 2023; Jiang et al., 2023; Meta, 2023), which are allegedly able to explain themselves in great detail and very convincingly (Chen et al., 2023; Agarwal et al., 2024).
Figure 5: Abstract diagram of the self-explanation paradigm, where the same model is trained to produce both the regular predictive output and an explanation, called a self-explanation. This paradigm is often seen with Large Language Models, where both the predictive output and the self-explanations appear as generated text.
Because the explanations are produced by a black-box this paradigm can be quite dangerous. Therefore, there must be solid evidence that the explanations are faithful for this approach to be valid. However, despite this immediate danger, the model that generates the explanation can in principle have access to all of the logic that produces the prediction. At a minimum, the same weights produce both the prediction and the explanation.
Importantly, self-explanations must relate to the model’s reasoning logic, not just the world or abstract concepts. However, presently there is little evidence that this is satisfied (Turpin et al., 2023; Lanham et al., 2023; Madsen et al., 2024a). This is not surprising, as the self-explanations are explicitly trained based on humans’ annotating how these explanations should look. However, humans don’t have any insight into how the model operates (Jacovi & Goldberg, 2020). As such, the model converges towards very convincing self-explanations with no regard for faithfulness (Agarwal et al., 2024; Chen et al., 2023).
While the other paradigms have found some solutions to their challenges, there are currently no known solutions to make self-explanations faithful. Even measuring faithfulness of self-explanation is very challenging (Huang et al., 2023). Currently there only exist a few metrics for specific self-explanations (Parcalabescu & Frank, 2023). However, future work may improve upon this, by developing more faithfulness metrics and by aligning models not just towards human preference but also faithfulness and truthfulness.
@@ 6 Limitations
This concern is orthogonal to the faithfulness concern and is therefore not discussed in this paper. However, readers are encouraged to study recent works like Schut et al. (2023); Kim (2022), which propose the new idea that it is not enough to frame explanations in terms that humans already understand. We should also develop new language and mental abstractions for humans to understand machines.
@@ 7 Conclusion
Although some evidence exists for the new paradigms presented in Section 5, these are, first and foremost, just ideas. It’s only in retrospect that we can truly know if one paradigm results in meaningful progress in the field. It is also entirely possible that neither of these ideas is what moves the interpretability field forward.
@@
Historically, a common tactic in post-hoc works was convincing visualization of explanations that aligned with our intuitions (Olah et al., 2017; Yosinski et al., 2015; Nguyen et al., 2016). However, such visualizations are empty arguments, as humans can’t know what a true explanation looks like (Geirhos et al., 2023). Likewise, intrinsic works have made seemingly strong theoretical arguments for why their methods are faithful, but these arguments failed to capture the whole model. Even the new learn-to-faithfully-explain paradigm have already shown sharp corners, where the explainer model unintentionally encode the prediction and is therefore unfaithful (Jethani et al., 2021).
@@
Finally, while these new paradigms are promising, it’s unlikely they will completely erase the current paradigms. We still teach both the particle and wave paradigms in physics. Most scientists don’t worry about whether there are true statements in math that cannot be proven.
@@
Acknowledgements
Sarath Chandar is supported by the Canada CIFAR AI Chairs program, the Canada Research Chair in Lifelong Machine Learning, and the NSERC Discovery Grant.
Siva Reddy is supported by the Facebook CIFAR AI Chairs program and NSERC Discovery Grant.